# Between Lines of Code: Unraveling the Distinct Patterns of Machine and Human Programmers

### 12 Jan 2024 v1

@@ 3.4.3. Token Frequency
Table 2 lists the top 50 tokens from human- and machine-authored code when $T$=0.2. Due to space limit, we omit the results when $T=1.0$, which has a similar result. From the results, we have several noteworthy observations:
Common Tokens: Human- and machine-authored code shares a commonality in their usage of certain tokens, including punctuation marks such as ‘.’, ‘(’, ‘)’, and structural keywords such as ‘if’, ‘return’, and ‘else’. This is because LLMs acquire foundational coding syntax after being trained on extensive human-written code corpora.
Error Handling: Tokens associated with error handling like ‘raise’ and ‘TypeError’, are more prevalent in machine-authored code. This difference implies that machine programmers emphasize robustness and exception handling more explicitly.
Programming Paradigms: Tokens indicating object-oriented programming like ‘self’ and ‘__init__’ are prominent in both human- and machine-authored code, which illustrates the model’s training alignment with this paradigm. However, machine-authored code appears to favor more boilerplate tokens like ‘__class__’ and also ‘__name__’, which could stem from its training on diverse object-oriented codebase.
Finding 3: Machine-authored code focuses more on exception handling and object-oriented principles than human, suggesting an emphasis on error prevention and adherence to common programming templates.
3.4.4. Token Category Proportion
Figure 3. Proportions of token categories in the studied code corpus
The analysis of token category proportions, as visualized in Figure 3, reveals intriguing insights into the coding conventions and stylistic nuances between human- and machine-authored code. The ‘keyword’, ‘operator’, ‘syntactic symbols’, and ‘whitespace’ proportions remain largely consistent between human and machine-authored code, across both temperatures. This consistency suggests that the foundational syntactical elements manifest similarly in both datasets.
Delving deeper into the more nuanced discrepancies, a few categories emerge that underscore the differential preferences or tendencies of human and machine writers:
@@ 3.4.4. Token Category Proportion
Finding 4: Machine-authored code tends to favor existing variables and data manipulation over new identifiers, and tends to have more comments to maintain clarity when the generation temperature grows.
3.4.5. Naturalness
Figure 4. Distribution of naturalness scores for human- and machine-authored code
Figure 4 shows that there is a great discrepancy of naturalness between machine- and human-authored code. Compared to human-authored code, the log likelihood scores of machine-authored code are mostly higher and the log rank scores are mostly lower, indicating that machine’s code is more “natural” than human-written code. Such observation in source code is consistent with the findings in natural language (Gehrmann et al., 2019; Ippolito et al., 2020; Mitchell et al., 2023; Solaiman et al., 2019).
Table 3. The naturalness of different token categories
@@ 3.4.5. Naturalness
Table 3 summarises the comparison results in terms of each token category at $T=0.2$. An intriguing finding is that whitespace tokens stand out with the highest deviation of naturalness, surpassing even the combined use of all tokens. This highlights a distinctive aspect of coding practices: machines, trained on extensive datasets, typically generate code with regular, predictable whitespace patterns. Humans, however, influenced by individual styles and practices, exhibit a wider variety in their use of whitespaces. The distinct patterns in machine-generated whitespaces, therefore, point to an inherent variation in coding style between machine and human.
@@
4. Detecting Machine Generated Code
The empirical results suggest that machine tends to write more concise and natural code with a narrower spectrum of tokens, towards regularization to programming principles, and the disparity is more pronounced in structural tokens such as whitespaces. This sparks a new idea for detecting machine-generated code: instead of perturbing arbitrary tokens, we focus on perturbing those structural tokens that best characterize the machine’s preference. Based on this idea, we introduce DetectCodeGPT, a novel zero-shot method for detecting machine-generated code. Our approach adapts the perturbation-based detection framework DetectGPT (Mitchell et al., 2023) to the code domain.
4.1. Problem Formulation
@@
4.2. Measuring Naturalness
Previous methods usually use the log likelihood of tokens to measure the naturalness of machine-authored content (Hindle et al., 2016; Solaiman et al., 2019). However, the log rank of tokens shows better performance comparing the naturalness of machine- and human-authored text (Mitchell et al., 2023; Gehrmann et al., 2019), because it offers a smoother and robust representation of token preference.
Unlike DetectGPT which directly calculates the log likelihood of tokens, we adopt the Normalized Perturbed Log Rank (NPR) score (Su et al., 2023) to capture the naturalness. The NPR score is formally defined as:
(3) $\mathbf{NPR}\left(x,p_{\theta},q\right)\triangleq\frac{\mathbb{E}_{\tilde{x}\sim q(\cdot\mid x)}\log r_{\theta}\left(\tilde{x}\right)}{\log r_{\theta}(x)},$
where $\log r_{\theta}(x)$ is the logarithm of the rank order of text $x$ sorted by likelihood under model $p_{\theta}$. In practice, $\mathbf{NPR}\left(x,p_{\theta},q\right)$ has been demonstrated to be more accurate for differentiating text origins, outperforming the log likelihood discrepancy $\mathbf{d}\left(x,p_{\theta},q\right)$ (Su et al., 2023).
4.3. Perturbation Strategy
Our empirical study indicates that the whitespace tokens serve as an important indicator of machine’s regularization and human’s diversity, which points to an inherent variation in structural coding style. Therefore, we propose an efficient and effective perturbation strategy with the following two types of perturbations below. Detailed explanations on the effectiveness of these perturbations are given in Section 6.1.
Space Insertion
(4) $n_{\text{spaces}}(c)\sim\mathcal{P}(\lambda_{\text{spaces}}).$
Newline Insertion
(5) $n_{\text{newlines}}(l)\sim\mathcal{P}(\lambda_{\text{newlines}}).$
We randomly choose one type of perturbation to the code snippet $x$ to generate a set of perturbed samples $\tilde{x}_{i}$ for $i\in[1,k]$, where $k$ is the number of perturbations. Through this step, we instill randomness at a granular structural level, thereby amplifying the perturbation’s efficacy. Our perturbation strategy introduces several distinct advantages over the conventional methods (Mitchell et al., 2023; Su et al., 2023) using MLM to perturb the code, which will be discussed in Section 6.2.
@@ 2       $p\sim\mathcal{U}(0,1)$;
             // Spaces Insertion for Perturbed Sample
@@ 8                   Insert $n_{\text{spaces}}(c)$ spaces at location $c$ in $x$;
             // Newlines Insertion for Perturbed Sample
@@ 22      Store the perturbed code as $\tilde{x}_{i}$;
25Calculate expectation: $\tilde{\mu}\leftarrow\frac{1}{k}\sum_{i}\mathbf{NPR}\left(\tilde{x}_{i},p_{\theta},q\right)$;
26 Estimate NPR for original code: $\text{NPR}_{x}\leftarrow\mathbf{NPR}\left(x,p_{\theta},q\right)-\tilde{\mu}$;
27 Calculate variance: $\tilde{\sigma}_{x}^{2}\leftarrow\frac{1}{k-1}\sum_{i}\left(\mathbf{NPR}\left(\tilde{x}_{i},p_{\theta},q\right)-\tilde{\mu}\right)^{2}$;
28 if $\frac{\text{NPR}_{x}}{\sqrt{\tilde{\sigma}_{x}}}>\epsilon$ then
@@
        // Probably machine-generated
29
30else
@@
        // Probably human-written
31
@@
Algorithm 1 DetectCodeGPT: Machine-Generated Code Detection with Structured Code Perturbation
Algorithm 1 summarizes the entire workflow of DetectCodeGPT. Our algorithm harnesses structured code perturbation to differentiate between human- and machine-authored code. At the core of our approach is the strategic insertion of spaces (Lines 5-9) and newlines (Lines 11-16) in code, a process that simulates the inherent randomness in human coding habits. The algorithm operates by generating perturbed versions of the code and then evaluating their NPR scores (Lines 20-22) with respect to the source model $\mathcal{M}$.
The threshold parameter $\epsilon$ in Line 23, pivotal for making the detection decision, offers flexibility in catering to different application scenarios. By adjusting $\epsilon$, users can balance between false positives and false negatives, tailoring the detection sensitivity according to the specific needs of the deployment context.
5. Evaluation
@@
5.1. Datasets
Apart from CodeSearchNet (Husain et al., 2020) described in Section 3.3, we select Python code from The Stack (Kocetkov et al., 2022) as another evaluation dataset. Similar to CodeSearchNet, The Stack provides codes from a variety of open-source projects representative of real-world scenarios. We use a parsed and filtered version (Manh et al., 2023) of this dataset and also concatenate the function definitions with their corresponding comments as prompts. For each combination of dataset and model, we construct 500 human and machine code pairs for evaluation. The maximum length of both human- and machine-authored code is trimmed to 128 tokens.
5.2. Evaluation Metric
Following prior works (Mitchell et al., 2023; Su et al., 2023), our primary metric for performance evaluation is the Area Under the Receiver Operating Characteristic curve (AUROC), which can be interpreted as the probability that a piece of machine-generated code is assigned a higher score than a piece of human-written code. Formally, given a set of true positive rates (TPR) and false positive rates (FPR) across different thresholds, the AUROC can be represented as:
@@ 5.3. Baselines and Models
  • Log $p(x)$ (Solaiman et al., 2019): Utilizes the source model’s average token-wise log probability to gauge code naturalness. Machine-generated code tends to have a higher score.
  • Entropy (Gehrmann et al., 2019): Interprets high average entropy in the model’s predictive distribution as indicative of machine generation.
  • (Log-) Rank (Gehrmann et al., 2019; Ippolito et al., 2020): The average observed rank or log rank of each token in the LLM prediction, with machine-generated passages typically showing smaller average values.
  • DetectGPT (Mitchell et al., 2023): Leverages the log probability of the original code and its perturbed variants to compute the perturbation discrepancy gap.
  • DetectLLM (Su et al., 2023): Introduces two methods, one blends log likelihood with log rank to compute LRR, and the other improves DetectGPT by incorporating the NPR score.
  • Supervised Method: a simple supervised detection method that predicts the authorship of a given code snippet using a binary classifier. Following OpenAI’s RoBERTa-based (Liu et al., 2019) GPT-2 detector [^2] https://github.com/openai/gpt-2-output-dataset/tree/master/detector, we employ CodeBERT (Feng et al., 2020) with an MLP head as the classifier and train it on a combination of 1000 samples generated by each model at each setting.
As for models to generate machine’s code, in consideration of GPU memory limitations, we judiciously selected a range of advanced LLMs with varying parameter counts, from 1 billion to 7 billion. We selected Incoder (Fried et al., 2022), Phi-1 (Gunasekar et al., 2023), StarCoder (Li et al., 2023), WizardCoder (Luo et al., 2023), CodeGen2 (Nijkamp et al., 2023) and CodeLlama (Rozière et al., 2023) from Huggingface  [^3] https://huggingface.co/models. This diverse array of models, each trained on distinct datasets and boasting different capacities, offers a comprehensive spectrum for evaluating our method.
5.4. Experimental Setup
For code generation with different models, we adopted the top-$p$ sampling strategy (Holtzman et al., 2020) with $p$=$0.95$ following (Chen et al., 2021). We explored two temperature settings, $T$=$0.2$ and $T$=$1.0$, as discussed in Section 3.2. The maximum length constraint for generated codes was set at $128$ tokens. With respect to the perturbation-specific hyperparameters, a grid search on a held-out set from the CodeSearchNet dataset, using the SantaCoder model (Allal et al., 2023), revealed optimal values. Consequently, we set $\alpha$ and $\beta$ to $0.5$, while $\lambda_{\text{spaces}}=3$ and $\lambda_{\text{newlines}}=2$. For all experiments, we maintained a consistent configuration of generating $50$ perturbations. For the DetectGPT and DetectLLM tasks, which involve an LLM in restoring perturbed code, we utilized the CodeT5+ (770M) model (Wang et al., 2023). All experiments are conducted on 2 NVIDIA RTX 4090 GPUs with 24GB memory.
Table 4. Performance (AUROC) of various detection methods
@@ 5.4. Experimental Setup
$\log p(x)$ Entropy Rank Log Rank  DetectGPT LRR NPR  Supervised  DetectCodeGPT
@@ 5.5. Detection Performance (RQ2)
Table 4 delineates the results of various methods. According to the results, DetectCodeGPT consistently outperforms baseline methods. In an impressive 21 of 24 combinations of dataset and model, our method provides the most accurate performance, which underscores its robustness across a variety of generative models, ranging from the 1.3 billion parameter InCoder to the 7 billion parameter CodeLlama. The high AUROC values achieved across these diverse settings confirm the method’s superior capability to generalize and reliably differentiate between machine-generated and human-written code.
@@
We can observe that the challenge of detection notably increases at a temperature setting of $T$=$1.0$ than $T$=$0.2$. This is possibly due to the higher randomness at this temperature, where models are likely to generate outputs with greater diversity. Despite these increased difficulties, the proposed method maintains its leading position in detection accuracy.
@@
It is also worth mentioning that our method, leveraging a zero-shot framework, frequently surpasses the accuracy of the supervised approach. The unstable performance of the supervised model reveals that it’s hard to detect machine-generated codes relying on the training data. The ability of our zero-shot approach to adapt and retain high performance under different levels of generative temperature variations is indicative of its sophisticated detection capability, making it a valuable tool for maintaining the integrity of codebases against the infiltration of machine-generated code in practical applications.
5.6. Ablation Study (RQ3)
Table 5. Performance of different perturbation strategies

### 24 Jan 2024 v2

@@ 3.4.1 Token Frequency
Table II lists the top 50 tokens from human- and machine-authored code when $T$=0.2. Due to space limit, we omit the results when $T$=1.0, which has a similar result. From the results, we have several noteworthy observations:
Common Tokens: Human- and machine-authored code shares a commonality in their usage of certain tokens, including punctuation marks such as “.”, “(”, “)”, and structural keywords such as “if”, “return”, and “else”. This is because LLMs acquire foundational coding syntax after being trained on extensive human-written code corpora.
Error Handling: Tokens associated with error handling like “raise” and “TypeError”, are more prevalent in machine-authored code. This difference implies that machine programmers emphasize robustness and exception handling more explicitly.
Programming Paradigms: Tokens indicating object-oriented programming like “self” and “__init__” are prominent in both human- and machine-authored code, which illustrates the model’s training alignment with this paradigm. However, machine-authored code appears to favor more boilerplate tokens like “__class__” and also “__name__”, which could stem from its training on diverse object-oriented codebase.
Finding 1: Machine-authored code focuses more on exception handling and object-oriented principles than human, suggesting an emphasis on error prevention and adherence to common programming templates.
3.4.2 Syntax Element Distribution
Figure 1: Syntax element distribution of the studied code corpus
The analysis of syntax element distributions, as visualized in Figure 1, reveals intriguing insights into the coding conventions and stylistic nuances between human- and machine-authored code. The “keyword”, “operator”, “syntactic symbols”, and “whitespace” proportions remain largely consistent between human and machine-authored code, across both temperatures. This consistency suggests that the foundational syntactical elements manifest similarly in both datasets.
Delving deeper into the more nuanced discrepancies, a few categories emerge that underscore the differential preferences or tendencies of human and machine writers (statistical significance $p<0.01$):
@@ 3.4.2 Syntax Element Distribution
Finding 2: Machine-authored code tends to favor existing variables and data manipulation over new identifiers, and tends to have more comments to maintain clarity when the generation temperature grows.
3.4.3 Zipf’s and Heaps’ Laws
(a) Zipf’s Law ($T=0.2$)
(b) Zipf’s Law ($T=1.0$)
(c) Heaps’ Law ($T=0.2$)
(d) Heaps’ Law ($T=1.0$)
Figure 2: Comparison of Zipf’s and Heaps’ laws on machine- and human-authored code
Finding 3: Machines demonstrate a preference for a limited spectrum of frequently-used tokens, whereas human code exhibits a richer diversity in token selection.
3.4.4 Number of Tokens and Lines
(a) Number of tokens ($T=0.2$)
(b) Number of tokens ($T=1.0$)
(c) Number of lines ($T=0.2$)
(d) Number of lines ($T=1.0$)
Figure 3: Distribution of code length for machine- and human-authored code
One could hypothesize several reasons for these observed patterns. The propensity for conciseness at lower temperatures may reflect LLM’s training data, where probably concise solutions were more prevalent or deemed more “correct”. On the flip side, human developers, often juggling multiple considerations like future code extensions, comments for peer developers, or even personal coding style, might craft lengthier solutions. Furthermore, the narrowing of disparities at higher temperatures can be attributed to the model’s increased willingness to explore varied coding styles. At higher temperatures, the LLM possibly mimics a broader spectrum of human coding patterns, capturing the essence of diverse coding habits and styles found in its training corpus.
Finding 4: Machines tend to write more concise code as instructed by their training objective, while human programmers tend to write longer code, reflective of their personal styles.
3.4.5 Token Likelihood and Rank
Figure 4: Distribution of naturalness scores for machine- and human-authored code
Figure 4 shows that there is a great discrepancy of naturalness between machine- and human-authored code. Compared to human-authored code, the log likelihood scores of machine-authored code are mostly higher and the log rank scores are mostly lower, indicating that machine’s code is more “natural” than human-written code. Such observation in source code is consistent with the findings in natural language [39, 38, 11, 40].
Table III: The naturalness of different categories of syntax elements. Statistical significance $p<0.001$.
@@ 3.4.5 Token Likelihood and Rank
Table III summarises the comparison results in terms of each token category at $T=0.2$. An intriguing finding is that whitespace tokens stand out with the highest deviation of naturalness, surpassing even the combined use of all tokens. This highlights a distinctive aspect of coding styles: machines, trained on extensive datasets, typically generate code with regular, predictable whitespace patterns. Humans, however, influenced by individual styles and practices, exhibit a wider variety in their use of whitespaces. The distinct patterns in machine-generated whitespaces, therefore, point to an inherent variation in coding style between machine and human.
@@
4 Detecting Machine Generated Code
The empirical results suggest that machine tends to write more concise and natural code with a narrower spectrum of tokens, towards regularization to programming principles, and the disparity is more pronounced in stylistic tokens such as whitespaces. This sparks a new idea for detecting machine-generated code: instead of perturbing arbitrary tokens, we focus on perturbing those stylistic tokens that best characterize the machine’s preference. Based on this idea, we introduce DetectCodeGPT, a novel zero-shot method for detecting machine-generated code. Our approach adapts the perturbation-based detection framework DetectGPT [11] to the code domain.
4.1 Problem Formulation
@@
4.2 Measuring Naturalness
Previous methods usually use the log likelihood of tokens to measure the naturalness of machine-authored content [32, 40]. However, the log rank of tokens shows better performance comparing the naturalness of machine- and human-authored text [11, 39], because it offers a smoother and robust representation of token preference.
Unlike DetectGPT which directly calculates the log likelihood of tokens, we adopt the Normalized Perturbed Log Rank (NPR) score [45] to capture the naturalness. The NPR score is formally defined as:
$\mathbf{NPR}\left(x,p_{\theta},q\right)\triangleq\frac{\mathbb{E}_{\tilde{x}\sim q(\cdot\mid x)}\log r_{\theta}\left(\tilde{x}\right)}{\log r_{\theta}(x)},$ (3)
where $\log r_{\theta}(x)$ is the logarithm of the rank order of text $x$ sorted by likelihood under model $p_{\theta}$. In practice, $\mathbf{NPR}\left(x,p_{\theta},q\right)$ has been demonstrated to be more accurate for differentiating text origins, outperforming the log likelihood discrepancy $\mathbf{d}\left(x,p_{\theta},q\right)$ [45].
@@ 2       $p\sim\mathcal{U}(0,1)$;
             // Spaces Insertion
@@ 8                   Insert $n_{\text{spaces}}(c)$ spaces at location $c$ in $x$;
             // Newlines Insertion
@@ 22      Store the perturbed code as $\tilde{x}_{i}$;
25Calculate expectation: $\tilde{\mu}\leftarrow\frac{1}{k}\sum_{i}\text{NPR}\left(\tilde{x}_{i},p_{\theta},q\right)$;
26 Estimate NPR for original code: $\text{NPR}_{x}\leftarrow\text{NPR}\left(x,p_{\theta},q\right)-\tilde{\mu}$;
[…]
@@
        // Probably machine-authored
28
29else
@@
        // Probably human-authored
30
@@
Algorithm 1 DetectCodeGPT: Machine-Generated Code Detection with Stylized Code Perturbation
4.3 Perturbation Strategy
Our empirical study indicates that the whitespace tokens serve as an important indicator of machine’s regularization and human’s diversity, which points to an inherent variation in coding style. Therefore, we propose an efficient and effective perturbation strategy with the following two types of perturbations below. Detailed explanations on the effectiveness of these perturbations are given in Section 6.1.
4.3.1 Space Insertion
$n_{\text{spaces}}(c)\sim\mathcal{P}(\lambda_{\text{spaces}}).$ (4)
4.3.2 Newline Insertion
$n_{\text{newlines}}(l)\sim\mathcal{P}(\lambda_{\text{newlines}}).$ (5)
We randomly choose one type of perturbation to the code snippet $x$ to generate a set of perturbed samples $\tilde{x}_{i}$ for $i\in[1,k]$, where $k$ is the number of perturbations. Through this step, we instill randomness at a granular stylistic level, thereby amplifying the perturbation’s efficacy. Our perturbation strategy introduces several distinct advantages over the conventional methods [11, 45] using MLM to perturb the code, which will be discussed in Section 6.2.
Algorithm 1 summarizes the entire workflow of DetectCodeGPT. Our algorithm harnesses stylized code perturbation to differentiate between human- and machine-authored code. At the core of our approach is the strategic insertion of spaces (Lines 4-9) and newlines (Lines 11-16) in code, a process that simulates the inherent randomness in human coding styles. The algorithm operates by generating perturbed versions of the code and then evaluating their NPR scores (Lines 20-21) with respect to the source model $\mathcal{M}$.
The threshold parameter $\epsilon$ in Line 22, pivotal for making the detection decision, offers flexibility in catering to different application scenarios. By adjusting $\epsilon$, users can balance between false positives and false negatives, tailoring the detection sensitivity according to the specific needs of the deployment context.
5 Evaluation
@@
5.1 Datasets
Apart from CodeSearchNet [21] described in Section 3.3, we select Python code from The Stack [22] as another evaluation dataset. Similar to CodeSearchNet, The Stack provides codes from a variety of open-source projects representative of real-world scenarios. We use a parsed and filtered version [46] of this dataset and also concatenate the function definitions with their corresponding comments as prompts as in [1]. For each combination of dataset and model, we construct 500 human and machine code pairs for evaluation. The maximum length of both human- and machine-authored code is trimmed to 128 tokens.
5.2 Evaluation Metric
Following prior works [11, 45], our primary metric for performance evaluation is the Area Under the Receiver Operating Characteristic curve (AUROC), which can be interpreted as the probability that a piece of machine-generated code is assigned a higher score than a piece of human-written code. Formally, given a set of true positive rates (TPR) and false positive rates (FPR) across different thresholds, the AUROC can be represented as:
@@ 5.3 Baselines and Models
  • Log $p(x)$ [40]: Utilizes the source model’s average token-wise log probability to gauge code naturalness. Machine-generated code tends to have a higher score.
  • Entropy [39]: Interprets high average entropy in the model’s predictive distribution as indicative of machine generation.
  • (Log-) Rank [39, 38]: The average observed rank or log rank of each token in the LLM prediction, with machine-generated passages typically showing smaller average values.
  • DetectGPT [11]: Leverages the log probability of the original code and its perturbed variants to compute the perturbation discrepancy gap.
  • DetectLLM [45]: Introduces two methods, one blends log likelihood with log rank to compute LRR, and the other improves DetectGPT by incorporating the NPR score.
  • GPTSniffer [47]: a supervised baseline that trains CodeBERT [48] to predict the authorship of a given code snippet. Following OpenAI’s RoBERTa-based [49] GPT detector [^2] https://github.com/openai/gpt-2-output-dataset/tree/master/detector, we train the model on a combination of 1000 samples generated by each model at each setting.
As for models to generate machine’s code, in consideration of GPU memory limitations, we judiciously selected a range of advanced LLMs with varying parameter counts, from 1 billion to 7 billion. We selected Incoder [24], Phi-1 [50], StarCoder [51], WizardCoder [25], CodeGen2 [4] and CodeLlama [6] from Huggingface  [^3] https://huggingface.co/models. This diverse array of models, each trained on distinct datasets and boasting different capacities, offers a comprehensive spectrum for evaluating our method.
5.4 Experimental Setup
For code generation with different models, we adopted the top-$p$ sampling strategy [41] with $p$=$0.95$ following [1]. We explored two temperature settings, $T$=$0.2$ and $T$=$1.0$, as discussed in Section 3.2. The maximum length constraint for generated codes was set at $128$ tokens. With respect to the perturbation-specific hyperparameters, a grid search on a held-out set from the CodeSearchNet dataset, using the SantaCoder model [3], revealed optimal values. Consequently, we set $\alpha$ and $\beta$ to $0.5$, while $\lambda_{\text{spaces}}=3$ and $\lambda_{\text{newlines}}=2$. For all experiments, we maintained a consistent configuration of generating $50$ perturbations. For the DetectGPT and DetectLLM, which involve an LLM in restoring perturbed code, we utilized the CodeT5+ (770M) model [52]. And as for the supervised baseline GPTSniffer, we trained the CodeBERT model for 5 epochs with a batch size of 16 and a learning rate of 2e-5 with AdamW optimizer. All experiments are conducted on 2 NVIDIA RTX 4090 GPUs with 24GB memory.
Table IV: Performance (AUROC) of various detection methods. Statistical significance $p<0.001$.
@@ 5.4 Experimental Setup
$\log p(x)$ Entropy Rank Log Rank  DetectGPT LRR NPR  GPTSniffer  DetectCodeGPT
@@ 5.5 Detection Performance (RQ1)
Table IV delineates the results of various methods. According to the results, DetectCodeGPT consistently outperforms baseline methods. Compared to the strongest baseline Log Rank, our method achieves an average relative improvement of 7.6% in AUROC. In an impressive 21 of 24 combinations of dataset and model, our method provides the most accurate performance, which underscores its robustness across a variety of generative models, ranging from the 1.3 billion parameter InCoder to the 7 billion parameter CodeLlama.
We also repeated the experiments 10 times and employed a paired t-test to assess the statistical significance of the performance differences between the methods, setting the significance level at 0.05. Results show that the performance superiority of our method was statistically significant, with p-values less than 0.001. The high AUROC scores achieved across these diverse settings confirm the method’s superior capability to generalize and reliably differentiate between machine-generated and human-written code.
@@
We can observe that the challenge of detection notably increases at a temperature setting of $T$=$1.0$ than $T$=$0.2$. This is possibly due to the higher randomness at this temperature, where models are likely to generate outputs with greater diversity in styles. Despite these increased difficulties, the proposed method maintains its leading position in detection accuracy.
@@
It is also worth mentioning that our method, leveraging a zero-shot framework, frequently surpasses the accuracy of the supervised GPTSniffer. The unstable performance of the supervised model reveals that it’s hard to detect machine-generated codes relying on the training data. The ability of our zero-shot approach to adapt and retain high performance under different levels of generative temperature variations is indicative of its sophisticated detection capability, making it a valuable tool for maintaining the integrity of codebases against the infiltration of machine-generated code in practical applications.
(a) Example 1
(b) Example 2
(c) Example 3
(d) Example 4
Figure 5: Examples of machine- and human-authored code snippets with corresponding predictions.
To further illustrate the effectiveness of our approach, we present some representative examples of machine-generated code detection by DetectCodeGPT and two most competitive baselines in Figure 5. We set the mean score of all the code snippets as the decision threshold $\epsilon$ for each method.
Examples 1 is a machine-generated code snippet. DetectCodeGPT can correctly detect it, while the baselines fail. In this example, the first and last “if” statements are separated from the rest of the code with newlines, and the two “if” statements in the middle are modularized together since they have similar functionalities. Such modularization and separation of code blocks are captured by DetectCodeGPT thanks to the stylized perturbation. Example 2 and 3 are human-written codes that are misclassified by other baselines, but DetectCodeGPT correctly identifies them. In Example 2, we can observe that the code blocks are sometimes separated with newlines (e.g., lines 5-7), but sometimes not (as seen with the rest of the code). In Example 3, although the code blocks are well separated with newlines, the human author omitted the spaces between operators “//” and “/”, but add spaces between “*” and “+”. Such freely and randomly stylized code reveals the inherent randomness in human coding habits. It’s hard for the baselines that only rely on token-wise conditional distributions to capture such randomness in coding styles. In contrast, DetectCodeGPT effectively leverages the style information to distinguish between machine-generated and human-written code again. These examples demonstrate the effectiveness of our method in detecting machine-generated code via stylized code perturbation.
However, there are also some cases where DetectCodeGPT fails to detect machine-generated code. Example 4 is a human-written code misclassified as machine-generated by all the approaches. We can observe that this code snippet is well-structured with newlines and spaces, and the comments are well-aligned with the code. Its resemblance to machine-generated code is striking, posing a significant challenge for distinction. This example highlights the difficulty of detecting machine-generated code among well-structured human-written codes with standard coding styles.
5.6 Ablation Study (RQ2)
Table V: Performance of different perturbation strategies